# A Training-Free Approach for Music Style Transfer with Latent Diffusion Models

### v1

@@ Abstract
Music style transfer, while offering exciting possibilities for personalized music generation, often requires extensive training or detailed textual descriptions. This paper introduces a novel training-free approach leveraging pre-trained Latent Diffusion Models (LDMs). By manipulating the self-attention features of the LDM, we effectively transfer the style of reference music onto content music without additional training. Our method achieves superior style transfer and melody preservation compared to existing methods. This work opens new creative avenues for personalized music generation.
@@ Introduction
Inspired by the success of neural style transfer in computer vision (Gatys 2015; Johnson, Alahi, and Fei-Fei 2016; Chandran et al. 2021; Kwon et al. 2024; Wang et al. 2023), music style transfer has emerged as a powerful technique for generating music by decomposing and recombining the content and style elements of different musical pieces (Dai, Zhang, and Xia 2018). Though the distinction between musical content and musical style is not formally defined and context-dependent, researchers have proposed various frameworks for decomposing musical elements into content and style components. Content typically encompasses structural elements such as melody, harmony, and rhythmic patterns, while style encompasses performance-specific attributes including timbre, articulation, dynamics, and genre-specific characteristics (Dai, Zhang, and Xia 2018; Cífka, Şimşekli, and Richard 2019). Several studies further demonstrated the possibility of style transfer using non-musical elements, such as environmental sound (Grinstein et al. 2018).
Recent advances in large language models (LLMs) and diffusion models have simplified music generation and style transfer, enabling intuitive control even for those without detailed musical knowledge. Conventional methods for diffusion model-based music generation utilize the generative power of pre-trained diffusion models. Agostinelli et al. (2023) and Copet et al. (2024) introduced music generation using text conditioning (text-to-audio generation) and melody guidance by using a transformer-based autoregressive model. These works make style transfer more simple and controllable by using text descriptions as style specifications and audio clips as content. Most recently, drawing inspiration from textual inversion in the diffusion model (Gal et al. 2022), Li et al. (2024) utilized a pseudo-word representation for musical styles, enabling style transfer for arbitrary input music while maintaining structural information (e.g., melody or rhythm). However, to obtain satisfactory style-transfer results, it is necessary to provide detailed textual guidance on various musical attributes (e.g., timbre, pitch, performance style, composition style), which require specialized knowledge. Additionally, training or fine-tuning these models demands substantial computational power and time, which greatly limits their practical application in real-world scenarios.
@@
To overcome these limitations, we propose a straightforward yet effective training-free approach for music style transfer using pre-trained latent diffusion models (LDM) (Rombach et al. 2022), which are commonly used for training text-to-image diffusion models. Building on the findings of (Chung, Hyun, and Heo 2024; Feng et al. 2022; Ma et al. 2024), attention maps are responsible for determining the spatial arrangement, while the key and value in cross-attention handle the content that fills the space. Additionally, (Chung, Hyun, and Heo 2024) showed that the self-attention layer in diffusion models is well-suited for style transfer, as it preserves the relationships between content image patches after the transfer and encourages style transfer based on the similarity of local textures between the content and the style. As a result, our method aims to transfer the style of reference music to the content music by explicitly manipulating the self-attention features of pre-trained large-scale diffusion models for text-to-image synthesis without any further training or optimization while leveraging the image characteristics of mel-spectograms. Specifically, we simply replace the content’s key and value of self-attention with those of the style music, focusing particularly on the later layers of the decoder that capture relevant local textures of mel-spectograms. To further improve music stylization, we also incorporate additional techniques originally introduced in (Chung, Hyun, and Heo 2024), including query preservation, attention temperature scaling, and initial latent Adaptive Instance Normalization (AdaIN).
@@
  • We propose a straightforward yet effective training-free approach for music style transfer, leveraging pre-trained LDM through direct manipulation of self-attention features.
  • Extensive experiments validate that our method can faithfully transfer the style of reference music to content music without the need for additional training or fine-tuning.
Figure 1: Overall architecture for our approach. Note that our approach does not require a training process
@@
Music Style Transfer. Music style transfer enables musical generation through element decomposition and recombination (Dai, Zhang, and Xia 2018). The field has evolved through various neural architectures, each addressing different aspects of musical transformation. Initial research focused on timbre transformation while preserving melodic content, with Engel et al. (2017) using WaveNet-based autoencoders and Grinstein et al. (2018) applying Convolutional Neural Networks (CNN) for timbre transfer between musical and non-musical sounds. These successes extended to genre transformation, where Brunner et al. (2018) demonstrated symbolic music manipulation using Recurrent Neural Networks (RNN). Cífka, Şimşekli, and Richard (2019) achieved more sophisticated genre transformation by separately processing melodic content and arrangement style by using the hybrid of CNN and RNN. Adversarial training approaches also produced significant results. WaveNet and CycleGAN architectures enabled style-preserved timbre transfer (Huang et al. 2018; Bonnici, Benning, and Saitis 2022), while Hung et al. (2019) achieved direct disentanglement of pitch and timbre. By leveraging adversarial approach, (Lee et al. 2020) demonstrated the potential of multi-media style transfer by injecting music as the style information for image style transfer.
Music style transfer has advanced significantly through both LLM and diffusion-based approaches. LLM-based methods, such as MusicLM (Agostinelli et al. 2023) and MusicGen (Copet et al. 2024), introduced transformer-based autoregressive models that enable text-conditioned and melody-guided generation, making style control more accessible through text descriptions and audio references. Huang et al. (2023) (Huang et al. 2023) first demonstrated diffusion models’ potential in music generation, working with both waveforms and spectrograms. Building on these advances and drawing inspiration from textual inversion image diffusion models (Gal et al. 2022), Li et al. (2024) (Li et al. 2024) developed a pseudo-word representation approach, achieving structure-preserving style transfer that maintains melodic and rhythmic elements with flexible style manipulation for arbitrary input music. However, all of these algorithms require detailed guidance or extensive model training. In this paper, we faithfully transfer the style to the content music without any optimization process.
@@
We utilized Stable Diffusion (Rombach et al. 2022) as the backbone to achieve music stylization, as shown in Figure  1. Our work is conducted in the image domain, processing a mel-spectogram obtained from the input audio waveform using a Short Time Fourier Transform (STFT), based on the ideas of Riffusion (Forsgren and Martiros 2022) and MusicTI (Li et al. 2024).
@@
Following (Chung, Hyun, and Heo 2024), we treat the features in the self-attention layers similarly to cross-attention, using the style image $I^{s}$ as the conditioning input. Specifically, during the generation process, we replace the key and value features of the content music’s mel-spectrogram with those from the style music’s mel-spectrogram. To achieve this, we first obtain the latent representations for both the content and style mel-spectograms through DDIM inversion (Song, Meng, and Ermon 2020). We then capture the self-attention features of the style mel-spectrogram throughout the DDIM inversion process. For pre-defined timesteps $t={0,...,T}$, we invert the style and content mel-spectrograms, denoted as $z_{0}^{m_{c}}$ and $z_{0}^{m_{s}}$, from the image space ($t=0$) to Gaussian noise ($t=T$). During this process, we also gather the query features of the content mel-spectrogram ($Q_{t}^{m_{c}}$) and the key and value features of the style mel-spectrogram ($K_{t}^{m_{s}}$, $V_{t}^{m_{s}}$) at each timestep. Next, we initialize the stylized output latent noise $z_{T}^{m_{o}}$ by directly copying the content latent noise $z_{T}^{m_{c}}$. The style transfer is carried out by replacing the original key $K_{t}^{m_{o}}$ and value $V_{t}^{m_{o}}$ in the self-attention layer with the key $K_{t}^{m_{s}}$ and value $V_{t}^{m_{s}}$ derived from the style mel-spectrogram, during the reverse process of generating the stylized output latent $z_{T}^{m_{o}}$. To preserve the content structure and avoid unwanted distortions, we implement a query preservation technique as follows:
$\begin{gathered}\bar{Q}_{t}^{m_{o}}=\alpha\times Q_{t}^{m_{c}}+(1-\alpha)\times Q_{t}^{m_{o}},\\ \phi_{out}^{m_{o}}=Attn({\bar{Q}_{t}^{m_{o}}},K_{t}^{m_{s}},V_{t}^{m_{s}}),\end{gathered}$ (3)
where $\alpha$ is a hyperparameter. These operations are applied to the later layers of the decoder (layers 7–12 in the SD model) that focus on capturing local texture features.
@@
Implementation Details
We experiment with the MusicTI Dataset (Li et al. 2024), which contains a total of 254 five-second clips, with 74 style clips and 179 content clips. Note that our approach does not require a training process; thus, we used all these data at inference only. For LDM, we used stable diffusion ver.1.5 (Rombach et al. 2022). In all our experiments, we fix the parameters of LDM and use the author-released codes using default configurations. All experiments were conducted using the PyTorch framework (Paszke et al. 2019) on a single NVIDIA A100(40G) GPU.
@@
In this section, the proposed model’s validity is assessed both qualitatively and quantitatively.
Figure 2 illustrates that our approach produces qualitatively satisfactory audio-transferred outputs throughout various instruments and musical elements. Analysis of mel-spectrograms shows that stylized outputs maintain content structure while incorporating characteristic elements of style sources. For instance, when applying accordion style to hip-hop audio, the stylized mel-spectrograms display distinctive horizontal energy bands in mid-frequencies, characteristic of accordion harmonics. Similarly, cornet-styled piano pieces exhibit concentrated mid-frequency energy and reduced high-frequency components, reflecting the cornet’s distinctive timbral characteristics. It is worth noting that even though the stable diffusion model is originally trained for text-to-image generation, not for music generation, it showed remarkable performance on synthesizing style-transferred music without any optimization process and music datasets.
content style
$FAD_{vgg}$ (↓) 9.46 24.49
$FAD_{CLAP}$ (↓) 1.10 1.23
Table 1: Quantitative results on FAD scores using pre-trained VGG and CLAP.
Conclusion & Future Works

### v2

@@ Abstract
Music style transfer enables personalized music creation by combining the structure of one piece with the stylistic characteristics of another. While recent approaches have explored text-conditioned generation and diffusion-based synthesis, most require extensive training, paired datasets, or detailed textual annotations. In this work, we introduce Stylus, a novel training-free framework for music style transfer that directly manipulates the self-attention layers of a pre-trained Latent Diffusion Model (LDM). Operating in the mel-spectrogram domain, Stylus transfers musical style by replacing key and value representations from the content audio with those of the style reference, without any fine-tuning. To enhance stylization quality and controllability, we further incorporate query preservation, CFG-inspired guidance scaling, multi-style interpolation, and phase-preserving reconstruction. Our method significantly improves perceptual quality and structural preservation compared to prior work, while remaining lightweight and easy to deploy. This work highlights the potential of diffusion-based attention manipulation for efficient, high-fidelity, and interpretable music generation—without training. Codes will be released upon acceptance.
@@ Introduction
Inspired by the success of neural style transfer in computer vision (Gatys 2015; Johnson, Alahi, and Fei-Fei 2016; Chandran et al. 2021; Kwon et al. 2024a; Wang et al. 2023), music style transfer has emerged as a powerful technique for generating music by decomposing and recombining the content and style elements of different musical pieces (Dai, Zhang, and Xia 2018). Though the distinction between musical content and musical style is not formally defined and context-dependent, researchers have proposed various frameworks for decomposing musical elements into content and style components. Content typically encompasses structural elements such as melody, harmony, and rhythmic patterns, while style encompasses performance-specific attributes including timbre, articulation, dynamics, and genre-specific characteristics (Dai, Zhang, and Xia 2018; C´ıfka, Şimşekli, and Richard 2019). Several studies further demonstrated the possibility of style transfer using non-musical elements, such as environmental sound (Grinstein et al. 2018). have enabled more intuitive control over music generation through explicit text descriptions. Significant advancements in text-conditioned music generation have emerged through diverse methodological approaches. Agostinelli et al. (2023) (Agostinelli et al. 2023) and Copet et al. (2024) (Copet et al. 2024) introduced transformer-based autoregressive models that enable music generation through text conditioning and melody guidance. In a complementary approach, Forsgren and Martiros (2022) (Forsgren and Martiros 2022) developed a text-to-music generation system based on denoised diffusion models operating on musical spectrograms. These methodologies collectively simplify style transfer by employing text descriptions as style specifications and audio clips as content sources, allowing users to effortlessly transform musical styles—particularly genres and moods—through textual prompts. However, to obtain satisfactory style-transfer results, it is necessary to provide detailed textual guidance on various musical attributes (e.g., timbre, pitch, performance style, and composition style), which require specialized knowledge. Furthermore, text-conditioned music generation or style transfer models have significant limitations in incorporating non-musical sound into their compositions, despite this being a widely adopted approach in contemporary music production.
Despite advances in text-based approaches, music-based music style transfer models have consistently demonstrated superior flexibility, expressivity, and creativity, leading to ongoing research in this domain. While early developments predominantly utilized Generative Adversarial Networks (GANs) for music-to-music style transfer and synthesis, recent research has shifted toward exploring Denoise Diffusion Models (DDMs), leveraging the generative capabilities of architectures originally pre-trained on natural images, such as Latent Diffusion Models (LDMs) (Rombach et al. 2022). A notable recent advancement by Li et al. (2024) (Li et al. 2024) draws inspiration from textual inversion techniques in diffusion models (Gal et al. 2022), employing pseudo-word representations for musical styles to enable style transfer for arbitrary input music while preserving structural elements like melody and rhythm. Despite these innovations, a significant limitation persists across both GAN-based and DDM-based approaches: the requirement to train or fine-tune style encoders for new applications. This necessity demands substantial computational resources and time, severely restricting the scalability and practical implementation of these models in real-world production environments.
@@
To overcome these limitations, we propose a straightforward yet effective training-free approach, Stylus, for music style transfer using pre-trained LDM, which are commonly used for training text-to-image diffusion models. Building on the findings of (Chung, Hyun, and Heo 2024; Feng et al. 2022; Ma et al. 2024), attention maps are responsible for determining the spatial arrangement, while the key and value in cross-attention handle the content that fills the space. Additionally, (Chung, Hyun, and Heo 2024) showed that the self-attention layer in diffusion models is well-suited for style transfer, as it preserves the relationships between content image patches after the transfer and encourages style transfer based on the similarity of local textures between the content and the style. As a result, our method aims to transfer the style of reference music to the content music by explicitly manipulating the self-attention features of pre-trained large-scale diffusion models for text-to-image synthesis without any further training or optimization while leveraging the image characteristics of mel-spectograms. Specifically, we simply replace the content’s key and value of self-attention with those of the style music, focusing particularly on the later layers of the decoder that capture relevant local textures of mel-spectograms. To enhance music stylization, we incorporate several key techniques. First, we adopt query preservation and initial latent Adaptive Instance Normalization (AdaIN) from (Chung, Hyun, and Heo 2024). We then introduce a novel method for controlling style injection, inspired by the success of Classifier-Free Guidance in image style transfer (Ho, Jain, and Abbeel 2020). Finally, we investigate a phase-preserving method for audio reconstruction from mel-spectrograms as an alternative to the standard Griffin-Lim algorithm(Griffin and Lim 1984) in music style transfer.
@@
  • We propose the first training-free music style transfer framework based on attention manipulation in a pre-trained diffusion model.
  • We introduce a suite of controllable and high-fidelity stylization techniques—including query preservation, guidance scaling, multi-style blending, and phase-preserving reconstruction.
  • We validate our approach through extensive experiments, showing that our method achieves strong results without any fine-tuning or optimization.
Figure 1: Overall architecture of Stylus. Note that our approach does not require any training process.
@@
Music Style Transfer. Music style transfer aims to generate music by decomposing and recombining musical elements (Dai, Zhang, and Xia 2018). Early studies focused on timbre transformation while preserving melodic content, using architectures such as WaveNet autoencoders (Engel et al. 2017) and CNNs (Grinstein et al. 2018). Subsequent works extended to genre transfer, leveraging symbolic representations (Brunner et al. 2018), hybrid CNN-RNN models (C´ıfka, Şimşekli, and Richard 2019), and adversarial frameworks including CycleGAN (Huang et al. 2018; Bonnici, Benning, and Saitis 2022) for timbre control and pitch disentanglement (Hung et al. 2019). Multi-modal extensions have also been explored (Lee et al. 2020).
Recent advances in large language models (LLMs) (Agostinelli et al. 2023; Copet et al. 2024; Kwon et al. 2024b) and diffusion-based methods (Huang et al. 2023) have enabled controllable music synthesis guided by text or brain-derived conditions. Building on this, Li et al. (Li et al. 2024) proposed a structure-preserving pseudo-token approach for flexible style manipulation. However, these methods typically rely on extensive training or conditioning. In contrast, our approach enables faithful, structure-aware style transfer without any further training or optimization process while achieving superior generation performance.
@@
We utilized Stable Diffusion (Rombach et al. 2022) as the backbone to achieve music stylization, as shown in Figure  1. Stylus is conducted in the image domain, processing a mel-spectogram obtained from the input audio waveform using a Short Time Fourier Transform (STFT), based on the conventions (Forsgren and Martiros 2022; Li et al. 2024).
@@
Following (Chung, Hyun, and Heo 2024), we treat the features in the self-attention layers similarly to cross-attention, using the style image $I^{s}$ as the conditioning input. Specifically, during the generation process, we replace the key and value features of the content music’s mel-spectrogram with those from the style music’s mel-spectrogram. To achieve this, we first obtain the latent representations for both the content and style mel-spectograms through DDIM inversion (Song, Meng, and Ermon 2020). We then capture the self-attention features of the style mel-spectrogram throughout the DDIM inversion process. For pre-defined timesteps t=0,,Tt={0,...,T}, we invert the style and content mel-spectrograms, denoted as $z_{0}^{m_{c}}$ and $z_{0}^{m_{s}}$, from the image space ($t=0$) to Gaussian noise ($t=T$). During this process, we also gather the query features of the content mel-spectrogram ($Q_{t}^{m_{c}}$) and the key and value features of the style mel-spectrogram ($K_{t}^{m_{s}}$, $V_{t}^{m_{s}}$) at each timestep. Next, we initialize the stylized output latent noise $z_{T}^{m_{o}}$ by directly copying the content latent noise $z_{T}^{m_{c}}$. The style transfer is carried out by replacing the original key $K_{t}^{m_{o}}$ and value $V_{t}^{m_{o}}$ in the self-attention layer with the key $K_{t}^{m_{s}}$ and value $V_{t}^{m_{s}}$ derived from the style mel-spectrogram, during the reverse process of generating the stylized output latent $z_{T}^{m_{o}}$. To preserve the content structure and avoid unwanted distortions, we implement a query preservation technique as follows:
$\begin{gathered}\bar{Q}_{t}^{m_{o}}=\gamma\times Q_{t}^{m_{c}}+(1-\gamma)\times Q_{t}^{m_{o}},\\ \phi_{out}^{m_{o}}=Attn({\bar{Q}_{t}^{m_{o}}},K_{t}^{m_{s}},V_{t}^{m_{s}}),\end{gathered}$ (3)
where $\gamma$ is a hyperparameter. These operations are applied to the later layers of the decoder (layers 7–12 in the SD model) that focus on capturing local texture features.
@@
Phase Preservation Strategy. While our model performs style transfer in the mel-spectrogram domain, we preserve the phase information directly from the original content audio. Instead of applying commonly used phase reconstruction methods such as Griffin-Lim, we reuse the phase extracted from the content audio via STFT and combine it with the stylized magnitude. This simple yet effective strategy significantly enhances the perceptual quality and temporal coherence of the generated output. We found that reusing the original phase not only reduces reconstruction artifacts but also better maintains the structural consistency of the source audio, despite the changes in timbre or style.
Controlling Stylistic Intensity via Guidance Scaling
To enable smooth and controllable music style transfer, we propose a training-free method that interpolates attention outputs inspired by Classifier-Free Guidance (CFG) (Ho, Jain, and Abbeel 2020).
While previous approaches achieve stylization by directly replacing the key and value components in self-attention layers, such binary replacement limits control over stylization strength and prevents seamless interpolation between multiple styles, especially in the music settings, where sophisticated continuous musical component like rhythms or melodies are important.
Instead of swapping key-value pairs, we compute attention outputs for both content and style representations independently. Given a query matrix $Q$ derived from a content mel-spectrogram and corresponding key-value pairs $(K_{c},V_{c})$ and $(K_{s},V_{s})$ from content and style respectively, we define the attention outputs as:
$\begin{gathered}\phi_{\text{content}}=\text{Attn}(Q,K_{c},V_{c}),\\ \phi_{\text{style}}=\text{Attn}(Q,K_{s},V_{s})\end{gathered}$ (5)
We then generate the guided attention output by linearly interpolating the two outputs:
$\begin{gathered}\phi_{\text{guided}}=\phi_{\text{content}}+\alpha\cdot(\phi_{\text{style}}-\phi_{\text{content}})\end{gathered}$ (6)
Here, $\alpha\in[0,1]$ is a style guidance scale that determines the degree of stylization. This formulation enables smooth control of stylization intensity and mitigates undesirable content distortion by maintaining proximity to the content representation (See Table  5).
Multi-style Interpolation. While guidance scaling allows fine-grained control over style intensity, our method can also be extended to support multi-style transfer. Given two style mel-spectrograms, we first compute their respective attention outputs $\phi_{\text{style}_{1}}$ and $\phi_{\text{style}_{2}}$, and interpolate them as follows:
$\begin{gathered}\phi_{\text{mix}}=(1-\lambda)\cdot\phi_{\text{style}_{1}}+\lambda\cdot\phi_{\text{style}_{2}},\\ \phi_{\text{guided}}=\phi_{\text{content}}+\beta\cdot(\phi_{\text{mix}}-\phi_{\text{content}})\end{gathered}$ (7)
where $\beta$ is a hyperparameter and $\lambda\in[0,1]$ controls the relative contribution of each style.
Implementation Details. We evaluate our method using the MusicTI dataset (Li et al. 2024), which contains 254 five-second audio clips, including 74 style clips and 179 content clips. Since our approach is entirely training-free, all data are used solely for inference. For the underlying latent diffusion model, we adopt the pretrained Stable Diffusion v1.5 (Rombach et al. 2022) with the default configuration and employ DDIM sampling (Song, Meng, and Ermon 2020) with 50 timesteps $(t=\{1,\ldots,50\})$. In our experiments, we use $\alpha=0.9$ and $\gamma=0.75$ as the default settings. All experiments are conducted using the PyTorch framework (Paszke et al. 2019) on a single NVIDIA A100(40G) GPU.
Table 1: Quantitative Comparison with various state-of-the art music generation algorithms. ”Musical Inst.” refers to musical instrument, and ”Pres.” denotes content preservation. The best results are set in bold. ↑: Higher is better. ↓: Lower is better.
Objective Evaluation User Study
Content Pres. ($\uparrow$) Musical Inst. Style Fit ($\uparrow$) Sound FX Style Fit ($\uparrow$) Content Pres. ($\uparrow$) Musical Inst. Style Fit ($\uparrow$) Sound FX Style Fit ($\uparrow$) Overall Quality ($\uparrow$)
MusicGen 0.312 0.218 0.014 2.71 2.57 1.94 2.64
MusicTI 0.396 0.263 0.118 3.11 3.24 3.21 2.79
Ours 0.531 0.162 0.132 4.28 3.23 3.26 3.41
Table 2: Efficiency Comparison with SOTA models aspect of inference time and GPU memory.
Inference Time (sec; $\downarrow$) Memory (GB; $\downarrow$)
MusicGen 13.86 16
MusicTI 11.78 8.7
Ours 10.16 18.1
Table 3: Ablation study of Style Intensity Sweeping (default: $\gamma$=0.75). $\alpha$ = 1 indicates only StyleID.
Content Musical Instrument Sound FX
Preservation ($\uparrow$) Style Fit ($\uparrow$) Style Fit ($\uparrow$)
$\alpha$ = 0 0.720 0.153 0.096
$\alpha$ = 0.25 0.678 0.154 0.101
$\alpha$ = 0.5 0.612 0.156 0.111
$\alpha$ = 0.75 0.535 0.157 0.128
$\alpha$ = 1 0.454 0.156 0.146
Table 4: Ablation study of Query Preservation Sweeping (default: $\alpha$=0.9).
Content Musical Instrument Sound FX
Preservation ($\uparrow$) Style Fit ($\uparrow$) Style Fit ($\uparrow$)
$\gamma$ = 0 0.242 0.144 0.161
$\gamma$ = 0.25 0.353 0.156 0.158
$\gamma$ = 0.5 0.435 0.157 0.149
$\gamma$ = 0.75 0.486 0.157 0.139
$\gamma$ = 1 0.519 0.158 0.133
Table 5: Ablation study of model (+ full replacement vs. CFG). Our results are calculated on the default setting ($\alpha$=0.9, $\gamma$=0.75). $\alpha$ = 1 indicates only StyleID, while $\alpha$ = 0 means CFG only.
Content Musical Instrument Sound FX
Preservation ($\uparrow$) Style Fit ($\uparrow$) Style Fit ($\uparrow$)
Ours 0.485 0.157 0.139
 - CFG (StyleID only) 0.454 0.156 0.146
 - Init AdaIN 0.520 0.161 0.133
 - Key features injection 0.209 0.088 0.134
 - Value features injection 0.152 0.082 0.117
 - phase preserving reconstruction 0.468 0.163 0.126
Table 6: Ablation study of Multi-style Interpolation ($\beta$; $\beta$ * style A + (1. - $\beta$) * style B).
Style Fit with Style A ($\uparrow$) Style Fit with Style B ($\uparrow$)
$\beta$ = 0.1 0.153 0.100
$\beta$ = 0.3 0.150 0.109
$\beta$ = 0.5 0.145 0.120
$\beta$ = 0.7 0.141 0.133
$\beta$ = 0.9 0.137 0.146
Table 7: Ablation study on different diffusion model versions, comparing between Stable Diffusion 1.5 and Stable Diffusion XL.
Content Musical Instrument Sound FX
Preservation ($\uparrow$) Style Fit ($\uparrow$) Style Fit ($\uparrow$)
SD v1.5 (Ours) 0.460 0.157 0.139
SDXL 0.122 0.013 -0.006
@@
In this section, the proposed model’s validity is assessed both qualitatively and quantitatively in comparison to state-of-the-art music style transfer approaches, including MusicTI (Li et al. 2024) and melody-guided MusicGen(Copet et al. 2024). For melody-guided MusicGen, we used content music as a guiding melody and transferred the style of content music with text descriptions of audio. Since the MusicTI requires training the style encoder, we trained MusicTI following the instructions reported in Li et al. 2024. The results for these methods are acquired by retraining the respective author-released codes using default configurations.
Qualitative Comparisons. Figure 2 illustrates that our approach produces qualitatively satisfactory audio-transferred outputs throughout various instruments and musical elements. Analysis of mel-spectrograms shows that stylized outputs maintain content structure while incorporating characteristic elements of style sources. Especially, our model superior to other models in robustly preserving structural elements of contents music, such as harmonics, melody and rhythm. For instance, when applying jaw harp style to relaxing music (used as content), our model successfully transfers the jaw harp’s periodic decay to the relaxing music by attenuating it’s high frequency energy. This precise replication of temporal dynamics effectively synthesizes a distinctive plucky rhythm and texture not present in the original. Similarly, our model preserves the musical structure of the ”color” music while successfully overlaying the ambient, noisy texture of the fire sound. Furthermore, our model achieves a successful timbral transformation by effectively erasing the clean, defined texture of the violin and replacing it with the characteristic breathy and indistinct texture of the cornet. Furthermore, our captures the accordion’s temporal characteristics by transforming the violin music’s short, detached notes into the long, sustained notes typical of an accordion performance, moving beyond simple timbral transfer. These results stand in contrast to other models that distorted the structure, highlighting our model’s ability to disentangle and reapply musical structure and non-musical texture.
Quantitative Comparisons. In line with previous studies on music style transfer  (Alinoori and Tzerpos 2022; C´ıfka et al. 2021; Li et al. 2024), we assess our model’s performance on content preservation and style fit. To evaluate content preservation, we calculate the CLAP score between the content and stylized mel-spectrograms. Style fit is measured using the CLAP score between the stylized output and a textual description of the style, which allows for a direct comparison with text-guided models, MusicGen-melody. We further analyze style fit across two distinct categories: musical instruments (e.g., accordion) and sound effects (e.g., fire), as they inherently have different properties regarding periodicity and harmonicity. Following Li et al. 2024, we established a reference score by computing the CLAP cosine similarity between textual descriptions and the original style mel-spectrograms. The average scores were 0.354 for musical instrument styles and 0.356 for sound effect styles. Of note, we evaluated our model on all 13,246 samples from the dataset established by Li et al. 2024, approximately four times the 3,222 samples used for previous models such as MusicGen and MusicTI. This comprehensive evaluation stems from our model’s key advantage: leveraging a pre-trained Stable Diffusion model (Rombach et al. 2022) enables one-shot style transfer without a dataset-specific style encoder, unlike methods that require training on multiple style audio samples. As shown in Table 1, our model significantly outperforms previous style transfer works in both content preservation and style fit for sound effects, validated by objective metrics.
User Study. Given the inherently subjective nature of evaluating stylization and audio fidelity, we conducted a user study to assess three key aspects: content preservation, stylization effectiveness, and overall audio quality. Participants were assigned 15 evaluation sets, each comprising a content, a style, and three randomly ordered stylized outputs generated by MusicGen, MusicTI, and Ours. No time constraints were imposed, allowing participants to independently assess each sample based on the aforementioned criteria. In total, we collected 1,386 individual responses, where each response corresponds to a single evaluation of one stylized output given a specific content-style pair. The user study results indicate that our method produces stylized audio with higher perceived quality compared to baseline approaches.
Ablation Studies
We also have conducted a series of ablation studies to provide justification for the architectural decisions employed and to highlight their effectiveness:
  • Style intensity control: Varying $\alpha$ from 0.0 to 1.0 demonstrates smooth transitions from content-preserving outputs to fully stylized results. A range of $\gamma$ values from 0.0 to 1.0 controls query preservation.
  • Architectural ablation: We isolate the role of key components in style injection: CFG, initial AdaIN layer, removing either key or value features, and phase preserving audio reconstruction from mel-spectogram images.
  • Multi-style Interpolation: We implemented vector linear interpolation of key and value features obtained from two different style audios’ mel-spectogram images.
In practice, we apply this blending method to decoder layers 7–12 of the Stable Diffusion U-Net, which are known to capture local texture features that reflect timbre and instrument characteristics in mel-spectrograms.
We randomly selected 30% of the samples from each category of both content and style audio, following prior research by evaluating the model on a subset of the data. The stratified sampling preserves diversity while reducing computational overhead.
First, we conducted an ablation study on two hyperparameters controlling style injection: $\gamma$ and $\alpha$. As shown in Tables 3 and 4, increasing $\alpha$ enhances style fit at the expense of content preservation, demonstrating a clear trade-off between the two. Interestingly, increasing $\gamma$, which preserves more of the content queries, consistently improves the style fit for musical instruments. We attribute this to the interaction between instrument timbre and temporally structured features (e.g., articulation, note transitions) maintained by content queries. In contrast, the style fit for sound effects degrades with higher $\gamma$, as their less-structured, event-based nature benefits more from style-dominated attention.
Second, we also conducted an architectural ablation study, with results presented in Table 5. The findings reveal a clear dichotomy in how components handle different styles. Removing Classifier-Free Guidance (CFG) degraded performance for musical instruments, suggesting CFG is crucial for injecting temporally structured stylistic features. In contrast, removing the initial AdaIN layer and phase matching hindered the stylization for sound effects, indicating its importance in transferring less-structured, event-based characteristics. Furthermore, excluding either key or value features during style injection critically degraded all metrics, confirming both are essential. These results demonstrate a trade-off in architectural components, where CFG is better suited for structured musical styles and the initial AdaIN excels with event-based sound effects. These findings reveal a trade-off in architectural components: CFG excels at handling structured musical styles, while the initial AdaIN is better suited for event-based sound effects.
Last, we investigated the capability of linear interpolation of our model. we observe that interpolating between two styles with different $\beta$ values enables continuous and perceptually smooth transitions. Although the perceptual quality of the mixture may vary depending on the compatibility of the styles, this approach offers a simple and flexible mechanism for exploring intermediate stylistic representations (see Table 7).
Discussions and Limitations
Our method enables training-free music style transfer with controllable style intensity and provides a comprehensive analysis of the architectural trade-offs between using musical instruments and sound effects as style references. A key limitation, however, is the model’s difficulty in distinguishing subtle timbral variations within the same instrument category. For instance, using two different jaw harp recordings as a style reference produces nearly identical outputs, failing to capture the nuanced differences a listener would perceive. Future music style transfer methods must address these fine-grained distinctions. To this end, our future work will focus on integrating models from computer vision adept at capturing subtle acoustic characteristics within mel-spectrograms. We believe this approach will significantly enhance the model’s ability to render nuanced and diverse stylistic variations.
Conclusion